Shift report 20020123 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 23 2002 - 15:48:38 EST

  • Next message: Flemming Videbaek: "Fw: SSH Closing"

    ------------------------------------------------------------
    Date: Wed Jan 23 15:48:37 2002, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: Flemming
    Log book#12 pages 107 - 111
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam for 11:20 - 13:43, about two hours.
       "Store lost due to power supply trip"
       "filling RHIC" at 15:35
    
    2. Pedestal and DVM : 6742, 6743, 6744
       
       At MRS@90(6B), FS@3(1/1A) 
          Run 6746, 6747 : for Trigger 4 and FS rate
                        roughly 10k trigger 4                  
          Run 6748  : MRS trigger only
          Run 6749  : only for 5min.
    
    3. No detector problems.
    ----------------------------
       T4,T5 tripped once. 
       T3 HV set default value.
       D3, D1, D2 tripped. --- Watch!!!
    
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Jan 23 2002 - 15:49:11 EST